OpenStudy (unicornpoopcookies):

URGENT PLEASE HELP use the distributive law to factor the following 2x+12+18y=

zepdrix (zepdrix):

Hmm, notice that the coefficients are all `even numbers`. So they're all divisible by 2, ya?

OpenStudy (unicornpoopcookies):

Yes I see that :)

OpenStudy (unicornpoopcookies):

12 divided by 2 6 and 18 divided by 2 9

OpenStudy (unicornpoopcookies):

6 and 9 divisible by 3

zepdrix (zepdrix):

So then,\[\large\rm 2x+12+18y=2\left(x+6+9y\right)\]I'm not sure how far they want us to factor this :3 Yah let's do that I suppose, take the 3 out of the 6 and 9.

OpenStudy (unicornpoopcookies):

I actually think thats how they want it but Im just having a hard time with the "laws"

OpenStudy (unicornpoopcookies):

Okay so from there we factor further right like to lowest terms

OpenStudy (unicornpoopcookies):

sorry my question mark dosent work :XXXXXX

zepdrix (zepdrix):

The distributive law looks like this:\[\large\rm a(b+c)=ab+ac\]I guess what we're doing in this problem is apply the distributive law `in reverse`.

zepdrix (zepdrix):

Yah let's try that. We'll take a 3 out of each of these blue terms. \[\large\rm 2\left(x+\color{royalblue}{6+9y}\right)=2\left(x+\color{royalblue}{3(2+3y)}\right)\]I divided 3 out of 6 to be left with a 2, and took a 3 out of 9 to be left with 3. ya?

OpenStudy (unicornpoopcookies):

yes

zepdrix (zepdrix):

I think that's all they wanted us to do :) lol weird problem.

OpenStudy (unicornpoopcookies):

writing this down thank you so much

zepdrix (zepdrix):

mmm that sparkly cookie looks delicious :D wait wait.. where did that come from :O GASP

OpenStudy (unicornpoopcookies):

Its a unicorn poop cookie

OpenStudy (unicornpoopcookies):

I know Ive been meaning to make some all these years since ive been on here lol

OpenStudy (unicornpoopcookies):

never have..

zepdrix (zepdrix):

\(\large \color{green}{(● ﹏☉)}\) ew